Diffusion Morning Report 07/03/06

Ted Berg tberg at snf.stanford.edu
Mon Jul 3 06:33:18 PDT 2006


Good Morning to the few who are here today,
    Looking good this morning so far. Epi HCl crossing over this 
morning. Tylan 1 is still down for TLC/O2 flow issues-investigating. LTO 
is coming
close to pull and clean.SEM seems OK for now. Hitachi Ebeam has some 
issues see report later. AMAT epi install moving ahead.  GeH4 install 
will be happening in about a week for as long as 2 weeks.  Everything 
else looks OK for now.

Diffusion Mike,Ray, Ted 	  	  	  	  	 
  	Date: 	7/3/06 	THICKNESS-Change at 	  	 
Tube Name/Number 	Status 	Comments 	  	  	 
Tylan 1 	Down 	flow issues 	  	  	 
Tylan 2 	Up 	  	  	  	 
Tylan 3 	Up 	  	  	  	 
Tylan 4 	Up 	  	  	  	 
Tylan 5-BBr3 	Up 	  	  	  	 
Tylan 6-POCl3 	Up 	  	  	  	 
Tylan 7-TEOS2 	Up 	  	1.8/30 	  	 
Tylan 8-SiGe 	Up 	  	33.96/30 	  	 
Tylan 9-Poly 	Up 	  	13.38/30 	  	 
Tylan 10-Nitride 	Up 	  	9.95/30 	  	 
Tylan 11-BPSG-LTO 	yellow 	  	28.26/30 	  	 
Tylan 12-FGA 	Up 	  	  	  	 
  	  	  	  	  	 
Thermco 1 	Up 	  	  	  	 
Thermco 2 	Up 	  	  	  	 
Thermco Poly 	Up 	  	  	  	 
Thermco Nitride 	Down 	Testing 	  	  	 
  	  	  	  	  	 
Epi 	Up 	  	  	  	 
  	  	  	  	  	 
Burn Box 	Up 	  	  	  	 
ZYGO 	up 	  	   	  	 
  	  	  	  	  	 
GAAS 22 	UP 	  	  	  	 
GAAS 23 	up 	  	  	  	 
  	  	  	  	  	 
SiH4 for furnaces 	  	A on line 360 psi 	 changeover working well  	  	 


-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20060703/de0d00ba/attachment.html>


More information about the diffusion mailing list